# Fusion of JWST data – Demonstrating practical feasibility

### 05 Mar 2026 v1

@@
Fusion of JWST data – Demonstrating practical feasibility
@@ 1 Introduction
Combining multiple data sets obtained with different instruments is a common practice in astrophysics. Usually, they are used independently to obtain complementary information on the chemical or physical characteristics of astronomical sources. Alternatively, one can directly combine some parts of the data sets, for instance as recently done by bacon2023muse who separated sources in MUSE datacubes by incorporating HST data. A similar but certainly more ambitious goal consists in performing a comprehensive data fusion, in which complementary observations are merged into a single data cube that retains the highest spatial and spectral resolutions of each sets (Fig. 1). While such a data fusion process has been widely used in Earth observation (yokoya2017hyperspectral), fusion of astronomical data has not yet been accomplished. This stems from several main challenges, notably that the astronomical wavelength ranges are typically large enough to result in non-negligible spectral variations of the optical point spread function (soulez2013restoration; hadj2017restoration), which significantly increases the complexity of the fusion process.
@@
Despite these challenges, some recent works have already investigated the fusion of synthetic astronomical data. Their main achievements have consisted in designing computationally efficient algorithms (guilloteau2020hyperspectral; pineau2023exact; pineau2025multispectral; lascar2025best) and generating realistic simulated data (guilloteau2020simulated). However, none of these previous studies considered real astronomical data. In this work, we demonstrate that – by leveraging the latest in-situ instrument models, the comprehensive NIRCam and NIRSpec documentation, and rigorous cross-calibration – a practical fusion of JWST data is now achievable.
@@ 2 Method
Figure 2: Selected JWST NIRCam images. a, Large field of view NIRCam image of the Orion Bar in the F210M filter. The red square denotes the field of view selected for the d203-506 protoplanetary disk. b, 1 x 1” images of the d203-506 protoplanetary disk observed by NIRCam filters F182M, F187N and F210M. c, 1.4 x 1.4” images of Titan observed by NIRCam filters F182M, F187N, F200W, F210M and F212N.
JWST conducts astronomical observations in the near- to mid-infrared range with an unprecedented quality (gardner2023james). Among its instruments, the Near Infrared Camera (NIRCam) (rieke2023NIRCamthroughput) imager and the Near Infrared Spectrograph (NIRSpec) integral field unit (IFU) (boker2022NIRSpecIFU) covers a spectral range from 0.6 to 5 $\mathrm{\mu m}$. NIRCam imager is composed of 29 filters and two acquisition channels: the short wavelength channel, below 2.35 $\mathrm{\mu m}$, that has a 0.031 arcsec pixel scale, and the long wavelength channel, above 2.35 $\mathrm{\mu m}$, that has a 0.063 arcsec pixel scale (STScINIRCamoverview). NIRSpec IFU has a 0.1 arcsec pixel scale and four high resolution filter/disperser combinations (STScINIRSpecIFUoverview). Its resolving power is around 2700, thus providing more than 9600 spectral channels.
We denote by $Y_{\mathrm{m}}$ the multispectral image acquired by the NIRCam imager and by $Y_{\mathrm{h}}$ the hyperspectral cube acquired by the NIRSpec IFU over the same scene. From this pair of data, we aim to reconstruct the underlying fused hyperspectral cube $X$ defined at NIRCam spatial resolution and NIRSpec spectral resolution. To solve this problem, we frame the fusion task as a regularised inverse problem which first requires to model the observational processes. The high spatial and spectral resolution cube $X$ to be reconstructed is assumed to be related to the NIRCam observations $Y_{\mathrm{m}}$ according to the forward model
@@
where $\textsf{NIRCam}(\cdot)$ stands for the operator accounting for the effects of NIRCam imager throughput and point spread function (PSF). Similarly for the NIRSpec IFU, the forward model writes
@@
where $\textsf{NIRSpec}(\cdot)$ resumes the spatial and spectral degradations resulting from the NIRSpec throughput, PSF and spatial subsampling. It is worth mentioning that for the two instruments the PSF are heavily varying across the wavelengths. These instrument models are detailed in Appendix A. Regarding the throughputs, the results reported hereafter are obtained by exploiting the latest available in-situ measurements (rieke2023NIRCamthroughput; giardino2022NIRSpecthroughput). We resorted to the theoretical PSF models derived by the WebbPSF simulation tool (perrin2014WebbPSF) for their balance of simplicity and accuracy over newer alternatives (nie2023hybpsf). Equations (1) and (2) relate the observation models to the observations through approximations. The expected errors underlying the approximation symbols $\approx$ are related to imperfect instruments models, data preprocessing, external and instrumental noises.
At this point, data fusion boils down to jointly inverting the two instrument models to recover the fused image $X$. However, despite its linearity, this inversion is ill-posed, making solutions non-unique and highly sensitive to model mismatches and noise. The proposed strategy consists in reformulating the fusion task as a regularised least squares problem
@@
where $\gamma$ is a parameter adjusting the importance of NIRSpec data fidelity term and $R(\cdot)$ is a regularisation. This formulation of the fusion task departs from the alternative that would exploit the noise statistics to derive a penalised log-likelihood. Indeed such an alternative requires to establish a proper noise model, which is very challenging, without ensuring substantial enhancements of the fusion results (pontoppidan2016pandeia; guilloteau2020simulated; guilloteau2020hyperspectral). Through regularisation, we promote expected spectral and spatial properties exhibited by the fused cube $X$. First, it is assumed to be linearly represented by a few elementary spectra. Hence, the adopted spectral regularisation imposes a low-rank structure on the data cube $X$, which is constrained to belong to an affine subspace of lower dimension. This spectral subspace is spanned by the most relevant spectra identified beforehand by a principal component analysis of the NIRSpec data. Then, the spatial regularisation is derived from a Sobolev norm to promote a smooth spatial content of the fused image. This choice has the advantage of leading to a globally quadratic minimisation problem, which can be easily solved (see Appendix B).
Figure 3: Selected JWST NIRSpec IFU F170LP filter and G235H disperser data. a, d203-506 protoplanetary disk observed by NIRSpec at 1.982 and 2.122 $\mathrm{\mu m}$ and b, spectra from those NIRSpec observations at the positions (red and black points) shown in the images. c, Titan observed by NIRSpec at 1.982 and 2.069 $\mathrm{\mu m}$ and d, spectra from those NIRSpec observations at the positions (red and black points) shown in the images. In the plot of the spectra, the two vertical dotted lines indicate the wavelengths of the images. Original data from the MAST database are rotated, aligned and cropped (we refer to this as co-registration) as described in Appendix A.
@@ 3 Application to JWST data
Our goal is to fuse pairs of input NIRSpec and NIRCam data that cover the same spectral range and field of view (FoV), a practical context herein referred to as symmetric fusion. Fusing real data present several key challenges: defining the feasible range for its application, finding corresponding data, and ensuring proper preprocessing such as alignment and cross-calibration. The methodology to meet these challenges is described in Appendix A. The proposed so-called Symmetric Fusion (SyFu) algorithm combines the preprocessing and fusion steps. In what follows, we present its application to JWST observations whose specifications are detailed in Appendix C.
@@ 3.1 JWST data
In principle, symmetric fusion of JWST data can be performed across six distinct wavelength ranges (see Fig. 5 and Table 3). Yet, we decided to restrict the available choices to the NIRCam short wavelength (SW) channel ($\lambda<2.35\penalty 10000\ \mathrm{\mu m}$) since it provides an angular resolution which is twice larger than the long wavelength channel. This choice thus allows the fusion process to reach the largest expected gain in terms of angular resolution. To ensure the largest spectral overlap between NIRCam SW and NIRSpec, the setup relies on the combination of the NIRSpec G235H disperser and F170LP filter (see Table 3). Five NIRCam filters cover this wavelength range and fully overlap with G235H/F170LP, namely the filters F182M, F187N, F200W, F210M and F212N (see Fig. 5 and Table 3).
Besides, ensuring that the pair of NIRCam and NIRSpec data to be fused covers the same FoV lead us to target JWST science programs with concomitant observations by the two instruments. Among the possible completed observations, we chose two different data sets available through the Mikulski Archive for Space Telescopes (MAST) that are part of two early release science (ERS) and guaranteed time observer (GTO) programs. More precisely, they correspond to the protoplanetary disk d203-506 in the Orion Bar observed within the PDRs4All ERS program (berne2022pdrs4all) and to Titan observed within the GTO program entitled “Titan climate, composition and clouds” (nixon2025atmosphere). These data sets have been elected to meet the aforementioned criteria defining symmetric fusion. The two pairs of NIRCam and NIRSpec JWST data sets are depicted in Fig. 2 and 3 and their respective properties are summarised in Table 2. In the case of d203-506, it is worth noting that only three filters (namely F182M, F187N and F210M) are available because no observations was planned through the filter F200W and those provided by the filter F212N were too noisy (habart2024pdrs4all). To mitigate the computational burden induced by the fusion algorithm, the two FoVs of interest where the fusion processes have been conducted are carefully limited to the celestial objects of interest. For the d203-506 data set, we chose a 1” $\times$ 1” FoV centred on the disk/jet (see Figures 2 and 3). For the Titan data set, we chose a 1.4” $\times$ 1.4” FoV covering the satellite (see Figures 2 and 3). Finally, we emphasise that the MAST archive may contain additional datasets that satisfy the criteria required for symmetric fusion.
@@ 3.2 Spatial and spectral inspections of the fused cubes
Figure 4: High resolution hyperspectral cubes resulting from JWST data fusion. a, d203-506 protoplanetary disk fused hyperspectral cube shown at 1.982 and 2.122 $\mathrm{\mu m}$ and b, spectra from this cube at the positions (red and black points) shown in the images. c, Titan fused hyperspectral cube shown at 1.982 and 2.069 $\mathrm{\mu m}$ and d, two spectra extracted from this cube at the positions (red and black points) shown in the images. In the plot of the spectra, the vertical dotted lines indicate the wavelengths of the images.
@@ 4 Discussion
The successful data fusion on real astronomical targets presented here opens up new possibilities for analysing astronomical data and designing future observations with JWST and beyond. In the case of JWST, the fused cubes achieve an angular resolution nearly three times higher than the native resolution of NIRSpec. This improvement could have significant implications across several fields of astronomy, as it may, for instance, enable to spatially resolve the thousands of emission lines present in the hundreds of irradiated protoplanetary disks in Orion (such as d203-506, presented in Fig. 4), offering new insights into their physical structure. These targets have recently garnered considerable interest (allen2025past; schroetter2025miri) and have already been observed with NIRCam (berne2022pdrs4all; mccaughrean2023jwst). Obtaining complementary NIRSpec IFU data would allow to fuse the data with NIRCam, ultimately yielding unique high-resolution data sets that could greatly enhance our understanding of planetary system formation, as the fusion could allow us to reach astronomical unit (AU) scales (the angular resolution of NIRCam corresponds to about 12 AU in Orion).
@@
There are also many promising opportunities for studying planet-forming disks closer to us, for which NIRCam data is already available—e.g., in Taurus L1527 (villenave2024jwst)—where the improved angular resolution from fusion could enable the identification of spatial structures within disks in specific emission lines.
@@
Nearby galaxies have also been observed with JWST, for instance as part of the PHANGS program with NIRCam (lee2023phangs). Targeted NIRSpec observations of specific regions could enable data fusion to study star-forming regions within these galaxies at unprecedented spatial scales and levels of detail. This could pave the way for fusion experiments for galaxies at greater distances. For instance, fusing NIRSpec and NIRCam observations of deep fields could help resolve individual star-forming regions in high-redshift galaxies (Claeyssens_2023), providing new tools to understand the formation of stellar clusters at cosmic dawn.
@@
A crucial requirement for accomplishing such a fusion is the availability, on the same telescope, of both an imager and spectrograph, as it is the case with NIRCam and NIRSpec on JWST. While not unique to JWST, this configuration is still relatively rare. For example, the X-IFU instrument on the future Athena mission could be combined with its imager (lascar2025best). We advocate for promoting coupled observation modes in future instruments, alongside with the development of dedicated data fusion algorithms integrated into their pipelines. This practice is already standard in the context of Earth observation.
@@
Although the successful data fusion presented here shows great promise, several improvements are needed to unlock its full potential. One limitation of the current SyFu algorithm is that both the imager and the IFU must observe the same FoV with complete spatial and spectral overlap (i.e., symmetric fusion). A way to expand the applicability of NIRCam–NIRSpec fusion would be to develop non-symmetric fusion methods. This would, for example, enable fusion across the entire NIRSpec wavelength range (0.6 to 5 $\mathrm{\mu m}$), despite the three wavelength gaps in the NIRCam data (Fig. 5).
@@
Also, in this work, we employed a classical regularisation based on the Sobolev norm. By design, this regularisation smooths the fused hyperspectral cube, which may lead to the loss of fine texture details present in the input image. A more advanced approach would involve weighting the Sobolev regularisation pixel-wise, for instance by exploiting the input high-resolution image to identify spatial contours that should be preserved, and then reducing the smoothing in those regions of the fused cube (guilloteau2022informed). Other alternatives include informing the regularisation using deep generative models, such as patch normalising flows (altekruger2023patchnr), to extract local structure from the high-resolution input image and improve the fusion process.
Appendix A Data preprocessing

### 27 Mar 2026 v2

# Fusion of JWST data: Demonstrating practical feasibility

@@
Fusion of JWST data: Demonstrating practical feasibility
@@ 1 Introduction
Combining multiple datasets obtained with different instruments is a common practice in astrophysics. Usually, different datasets are used independently to obtain complementary information on the chemical or physical characteristics of astronomical sources. Alternatively, one can directly combine some parts of datasets. For instance, bacon2023muse separated sources in MUSE data cubes by incorporating Hubble Space Telescope data. A similar but certainly more ambitious goal consists of performing a comprehensive data fusion in which complementary observations are merged into a single data cube that retains the highest spatial and spectral resolutions of each set (Fig. 1). While such a data fusion process has been widely used in Earth observation (yokoya2017hyperspectral), fusion of astronomical data has not yet been accomplished due to several main challenges, notably that the astronomical wavelength ranges are typically large enough to result in non-negligible spectral variations of the optical point spread function (PSF) (soulez2013restoration; hadj2017restoration), which significantly increases the complexity of the fusion process.
@@
Despite these challenges, some recent works have already investigated the fusion of synthetic astronomical data. Their main achievements include designing computationally efficient algorithms (guilloteau2020hyperspectral; pineau2023exact; pineau2025multispectral; lascar2025best) and generating realistic simulated data (guilloteau2020simulated). However, none of the previous studies has considered real astronomical data. In this work, we demonstrate that—by leveraging the latest in situ instrument models, the comprehensive Near Infrared Camera (NIRCam) and Near Infrared Spectrograph (NIRSpec) documentation, and rigorous cross-calibration—a practical fusion of JWST data is now achievable.
@@ 2 Method
Figure 2: Selected JWST NIRCam images. Large field of view NIRCam image (a) of the Orion Bar in the F210M filter. The red square denotes the field of view selected for the d203-506 protoplanetary disk. 1 x 1” images (b) of the d203-506 protoplanetary disk observed by NIRCam filters F182M, F187N and F210M. 1.4 x 1.4” images c of Titan observed by NIRCam filters F182M, F187N, F200W, F210M, and F212N.
The JWST conducts astronomical observations in the near- to mid-infrared range with unprecedented quality (gardner2023james). Among its instruments, the NIRCam (rieke2023NIRCamthroughput) imager and the NIRSpec integral field unit (IFU) (boker2022NIRSpecIFU) covers a spectral range from 0.6 to 5 $\mathrm{\mu m}$. The NIRCam imager is composed of 29 filters and two acquisition channels: the short wavelength channel, below 2.35 $\mathrm{\mu m}$, which has a 0.031 arcsec pixel scale, and the long wavelength channel, above 2.35 $\mathrm{\mu m}$, which has a 0.063 arcsec pixel scale (STScINIRCamoverview). The NIRSpec IFU has a 0.1 arcsec pixel scale and four high resolution filter and disperser combinations (STScINIRSpecIFUoverview). Its resolving power is around 2700, thus providing more than 9600 spectral channels.
We denote by $Y_{\mathrm{m}}$ the multispectral image acquired by the NIRCam imager and by $Y_{\mathrm{h}}$ the hyperspectral cube acquired by the NIRSpec IFU over the same scene. From this pair of data, we aim to reconstruct the underlying fused hyperspectral cube $X$ defined at the NIRCam spatial resolution and the NIRSpec spectral resolution. To solve this problem, we framed the fusion task as a regularised inverse problem that first requires modelling of the observational processes. The high spatial and spectral resolution cube ($X$) to be reconstructed is assumed to be related to the NIRCam observations ($Y_{\mathrm{m}}$) according to the forward model
@@
where $\textsf{NIRCam}(\cdot)$ stands for the operator accounting for the effects of the NIRCam imager throughput and PSF. Similarly for the NIRSpec IFU, the forward model is written as
@@
where $\textsf{NIRSpec}(\cdot)$ resumes the spatial and spectral degradations resulting from the NIRSpec throughput, PSF, and spatial sub-sampling. It is worth mentioning that for the two instruments the PSF varies heavily across the wavelengths. The instrument models are detailed in Appendix B. Regarding the throughputs, the results reported hereafter were obtained by exploiting the latest available in situ measurements (rieke2023NIRCamthroughput; giardino2022NIRSpecthroughput). We resorted to the theoretical PSF models derived by the WebbPSF simulation tool (perrin2014WebbPSF) for their balance of simplicity and accuracy over newer alternatives (nie2023hybpsf). Equations (1) and (2) relate the observation models to the observations through approximations. The expected errors underlying the approximation symbols ($\approx$) are related to imperfect instruments models, data pre-processing, and external and instrumental noises.
At this point, data fusion boils down to jointly inverting the two instrument models to recover the fused image, $X$. However, despite its linearity, this inversion is ill-posed, making solutions non-unique and highly sensitive to model mismatches and noise. The strategy we propose consists of reformulating the fusion task as a regularised least squares problem:
@@
where $\gamma$ is a parameter adjusting the importance of the NIRSpec data fidelity term and $R(\cdot)$ is a regularisation. This formulation of the fusion task departs from the alternative that would exploit the noise statistics to derive a penalised log-likelihood. Indeed such an alternative requires establishing a proper noise model, which is very challenging without ensuring substantial enhancements of the fusion results (pontoppidan2016pandeia; guilloteau2020simulated; guilloteau2020hyperspectral). Through regularisation, we promote expected spectral and spatial properties exhibited by the fused cube $X$. First, $X$ is assumed to be linearly represented by a few elementary spectra. Hence, the adopted spectral regularisation imposes a low-rank structure on the data cube $X$, which is constrained to belong to an affine subspace of lower dimension. This spectral subspace is spanned by the most relevant spectra identified beforehand by a principal component analysis of the NIRSpec data. Then, the spatial regularisation is derived from a Sobolev norm to promote a smooth spatial content of the fused image. This choice has the advantage of leading to a globally quadratic minimisation problem, which can be easily solved (see Appendix C).
Figure 3: Selected JWST NIRSpec IFU F170LP filter and G235H disperser data. NIRSpec observations of the d203-506 protoplanetary disk at 1.982 and 2.122 $\mathrm{\mu m}$ (a). Spectra from those observations (b) at the positions (red and black points) shown in the images. Titan observed by NIRSpec (c) at 1.982 and 2.069 $\mathrm{\mu m}$. Spectra from those observations (d) at the positions (red and black points) shown in the images. In the plot of the spectra, the two vertical dotted lines indicate the wavelengths of the images. Original data from the MAST database have been rotated, aligned, and cropped (we refer to this as co-registration) as described in Appendix B.
@@ 3 Application to JWST data
Our goal is to fuse pairs of input NIRSpec and NIRCam data that cover the same spectral range and field of view (FoV), a practical context herein referred to as symmetric fusion. Fusing real data presents several key challenges: defining the feasible range for its application, finding corresponding data, and ensuring proper pre-processing, such as alignment and cross-calibration. The methodology to meet these challenges is described in Appendix B. The proposed so-called Symmetric Fusion (SyFu) algorithm combines the pre-processing and fusion steps. In what follows, we present its application to JWST observations whose specifications are detailed in Appendix D.
@@ 3.1 JWST data
In principle, symmetric fusion of JWST data can be performed across six distinct wavelength ranges (see Fig. 5 and Table 3). Yet we decided to restrict the available choices to the NIRCam short wavelength channel ($\lambda<2.35\penalty 10000\ \mathrm{\mu m}$) since it provides an angular resolution that is two times higher than the long wavelength channel. This choice thus allows the fusion process to reach the largest expected gain in terms of angular resolution. To ensure the largest spectral overlap between the NIRCam short wavelength channel and NIRSpec, the setup relies on the combination of the NIRSpec G235H disperser and F170LP filter (see Table 3). Five NIRCam filters cover this wavelength range and fully overlap with G235H/F170LP. Namely, the filters are F182M, F187N, F200W, F210M, and F212N (see Fig. 5 and Table 3).
In addition, ensuring that the pair of NIRCam and NIRSpec data to be fused cover the same FoV led us to target JWST science programs with concomitant observations by the two instruments. Among the possible completed observations, we chose two different datasets available through the Mikulski Archive for Space Telescopes (MAST) that are part of two Early Release Science (ERS) and Guaranteed Time Observation (GTO) programs. More precisely, they correspond to the protoplanetary disk d203-506 in the Orion Bar observed within the PDRs4All ERS program (berne2022pdrs4all) and to Titan observed within the GTO program entitled ‘Titan climate, composition and clouds’ (nixon2025atmosphere). These datasets have been elected to meet the aforementioned criteria defining symmetric fusion. The two pairs of NIRCam and NIRSpec JWST datasets are depicted in Fig. 2 and 3, and their respective properties are summarised in Table 2. In the case of d203-506, it is worth noting that only three filters (namely F182M, F187N, and F210M) are available because no observations were planned through the filter F200W, and those provided by the filter F212N were too noisy (habart2024pdrs4all). To mitigate the computational burden induced by the fusion algorithm, the two FoVs of interest where the fusion processes have been conducted were carefully limited to the celestial objects of interest. For the d203-506 dataset, we chose a 1” $\times$ 1” FoV centred on the disk or jet (see Figures 2 and 3). For the Titan dataset, we chose a 1.4” $\times$ 1.4” FoV covering the satellite (see Figures 2 and 3). Finally, we emphasise that the MAST archive may contain additional datasets that satisfy the criteria required for symmetric fusion.
@@ 3.2 Spatial and spectral inspections of the fused cubes
Figure 4: High resolution hyperspectral cubes resulting from JWST data fusion. The d203-506 protoplanetary disk fused hyperspectral cube (a) shown at 1.982 and 2.122 $\mathrm{\mu m}$. Spectra from this cube (b) at the positions (red and black points) shown in the images. Titan fused hyperspectral cube (c) shown at 1.982 and 2.069 $\mathrm{\mu m}$. Two spectra extracted from this cube (d) at the positions (red and black points) shown in the images. In the plot of the spectra, the vertical dotted lines indicate the wavelengths of the images.
@@ 4 Discussion
The successful data fusion on real astronomical targets presented here opens up new possibilities for analysing astronomical data and designing future observations with JWST and beyond. In the case of JWST, the fused cubes achieve an angular resolution nearly three times higher than the native resolution of NIRSpec. This improvement could have significant implications across several fields of astronomy, as it may, for instance, enable spatially resolving the thousands of emission lines present in the hundreds of irradiated protoplanetary disks in Orion (such as d203-506, presented in Fig. 4), which would offer new insights into their physical structure. These targets have recently garnered considerable interest (allen2025past; schroetter2025miri) and have already been observed with NIRCam (berne2022pdrs4all; mccaughrean2023jwst). Obtaining complementary NIRSpec IFU data would allow a fusion with NIRCam data and ultimately yield unique high-resolution datasets that could greatly enhance our understanding of planetary system formation, as the fusion would allow us to reach astronomical unit scales (the angular resolution of NIRCam corresponds to about 12 AU in Orion).
@@
There are also many promising opportunities for studying planet-forming disks closer to us and for which NIRCam data are already available (e.g. in Taurus L1527 (villenave2024jwst)). In such cases, the improved angular resolution from fusion could enable the identification of spatial structures within disks in specific emission lines.
@@
Nearby galaxies have also been observed with JWST, for instance, as part of the PHANGS program with NIRCam (lee2023phangs). Targeted NIRSpec observations of specific regions could enable data fusion to study star-forming regions within these galaxies at unprecedented spatial scales and levels of detail. This could pave the way for fusion experiments on galaxies at greater distances. For instance, fusing NIRSpec and NIRCam observations of deep fields could help resolve individual star-forming regions in high-redshift galaxies (Claeyssens_2023), providing new tools to understand the formation of stellar clusters at cosmic dawn.
@@
A crucial requirement for accomplishing such a fusion is the availability, on the same telescope, of both an imager and spectrograph, as is the case with NIRCam and NIRSpec on JWST. While not unique to JWST, this configuration is still relatively rare. For example, the X-IFU instrument on the future Athena mission could be combined with its imager (lascar2025best). We advocate for promoting coupled observation modes in future instruments alongside the development of dedicated data fusion algorithms integrated into their pipelines. This practice is already standard in the context of Earth observation.
@@
Although the successful data fusion presented here shows great promise, several improvements are needed to unlock its full potential. One limitation of the current SyFu algorithm is that both the imager and the IFU must observe the same FoV with complete spatial and spectral overlap (i.e. symmetric fusion). A way to expand the applicability of NIRCam–NIRSpec fusion would be to develop non-symmetric fusion methods. This would, for example, enable fusion across the entire NIRSpec wavelength range (0.6 to 5 $\mathrm{\mu m}$), despite the three wavelength gaps in the NIRCam data (Fig. 5).
@@
In this work, we also employed a classical regularisation based on the Sobolev norm. By design, this regularisation smooths the fused hyperspectral cube, which may lead to the loss of fine texture details present in the input image. A more advanced approach would involve weighting the Sobolev regularisation pixel-wise, for instance, by exploiting the input high-resolution image to identify spatial contours that should be preserved and then reducing the smoothing in those regions of the fused cube (guilloteau2022informed). Other alternatives include informing the regularisation using deep generative models, such as patch normalising flows (altekruger2023patchnr), to extract local structure from the high-resolution input image and improve the fusion process.
Appendix A Data availability
Appendix B Data pre-processing